OpenStudy (anonymous):

What is the sum of all even integers between 99 and 301?

OpenStudy (anonymous):

first term = 100 , last term= 300 , common difference =2 no. of terms = (300-100)/2 = 100 Now, sum of all even numbers between 99 and 301 = 100/2 * (100 + 300) = 50*400=20000

OpenStudy (amistre64):

would of liked to have seen some leeway for the asker to participate in the process as opposed to just having the whole thing solved out without them having to even bat a synapse ....

OpenStudy (amistre64):

would *have ? would *of? .... think its have :)

OpenStudy (uri):

have..

OpenStudy (debbieg):

"would have".... and I couldn't agree more.

OpenStudy (amistre64):

if staying around to teach the process is a burden, then we can at least post the process and let them determine the final part or ask questions about what they do not understand.

OpenStudy (uri):

I remember a girl asking for ansurs yesterday reported her many times...and i saw mods coming and seeing..still she was online and posting questions...this happened 3 times.

OpenStudy (debbieg):

And the worst, like yesterday, is when I spent 10-15 minutes trying to help lead someone through that process, then someone else pops on the thread and spits out the answer, and the original asker medals them and is done! grrrr.

OpenStudy (amistre64):

they get warned .... and if the behaviour is blantant enough they get a suspension of a limited duration

OpenStudy (amistre64):

yeah :) thats defined as sniping in the code of conduct; and its blatant enough to determine a suspension for the asker

OpenStudy (debbieg):

Oh good. I'm new here so still learning the ropes. :)

OpenStudy (uri):

@DebbieG Ikr? that pisses me off.

OpenStudy (amistre64):

what gets me is when trying to lure a response from a poster you get: whats the answer? so whats the final answer? idk .. etc

OpenStudy (uri):

Lol..give wrong ansurs duh..